Live GMAT classes with Target Test Prep start Nov 21 and Nov 22. First class free.

See classes
Target Test Prep · GMAT

Choose how you want to prepare

Learn live with an expert or move at your own pace. Every option includes the complete TTP study system.

Newsweek Readers’ Choice Awards 2026 winner: Target Test Prep, #1 Test Prep

Target Test Prep was voted #1 Best Test Prep in the 2026 Newsweek Readers’ Choice™ Awards.

★★★★★5.0563 reviews
GMATLiveTeachStarts Nov 22
Josh Braslow, Target Test Prep GMAT expert
LIVE ONLINE CLASSES

Get Ready for GMAT Test Day Faster with Live Online Classes

with Josh Braslow, 20+ years in test prep

Nov 22 · Josh Braslow
Sun · 9:30 AM to 12:30 PM ET
Included
40 hours of live online classes + 6 months of TTP OnDemand
  • Attend the first class for free
  • Every class is recorded, so you never fall behind
View classes & enroll
Limited seats availableTarget Test Prep
EALiveTeachOnDemandStart anytime
EXECUTIVE ASSESSMENT

Target Test Prep EA OnDemand

Self-paced EA prep. Study on your schedule.

EXECUTIVE ASSESSMENT

with

165+ EA score guarantee
$05-day trial no automatic billing
Schedule
Included
40 hours of live online classes plus six months of access to the complete TTP EA OnDemand course.
  • 165+ EA Score Guarantee
  • 4,100+ Quant, Verbal, and Integrated Reasoning practice questions
  • 400+ hours of in-depth video lessons
  • 3,000+ step-by-step video solutions
View EA class & enroll Start free 5-day trial
Limited cohort · enrollment openTrial includes full course accessTarget Test Prep
GMATOnDemandStart anytime
SELF-PACED MASTERCLASS

Target Test Prep GMAT OnDemand

Complete access from day one. Study on your schedule.

715+ score guarantee
$0to start then $127/mo
  • Personalized study plan and analytics
  • Thousands of lessons and practice questions

If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one

Expert replies
by BTGmoderatorDC » Mon Apr 06, 2020 5:35 pm
Timer
00:00
Your answer
Difficulty–

right wrong

Pick an answer to see how members did

If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one tail?

A) 3/4
B) 11/12
C) 15/16
D) 31/32
E) 63/64


OA C

Source: Veritas Prep
Join the discussion
Source: — Problem Solving |

BTGmoderatorDC wrote: ↑
Mon Apr 06, 2020 5:35 pm
If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one tail?

A) 3/4
B) 11/12
C) 15/16
D) 31/32
E) 63/64

OA C

Source: Veritas Prep
For a fair coin, the probability of getting a head or a tail = 1/2

She can get at least one head and one tail in the following ways: 1 head, 4 tails; 2 heads, 3 tails;3 heads, 2 tails;4 heads, 1 tail.

• Probability of getting 1 head and 4 tails = 5C1*(1/2)*(1/2)^4 = 5/32;
• Probability of getting 2 heads and 3 tails = 5C2*(1/2)^2*(1/2)^3 = 10/32;
• Probability of getting 3 heads and 2 tails = 5C3*(1/2)^3*(1/2)^2 = 10/32;
• Probability of getting 4 heads and 1 tail = 5C4*(1/2)^4*(1/2) = 5/32

So, the probability that she will receive at least one head and one tail
= 5/32 + 10/32 + 10/32 + 5/32 = 30/32 = 15/16

The correct answer: C

Hope this helps!

-Jay
_________________
Manhattan Review Test Prep

Locations: Manhattan Review Bangalore | Free GRE Practice Test | GRE Prep Hyderabad | Kukatpally GRE Coaching | and many more...

Schedule your free consultation with an experienced GMAT Prep Advisor! Click here.
Join the discussion

BTGmoderatorDC wrote: ↑
Mon Apr 06, 2020 5:35 pm
If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one tail?

A) 3/4
B) 11/12
C) 15/16
D) 31/32
E) 63/64


OA C

Source: Veritas Prep
If a coin is flipped 5 times in a row, there are 32 possible outcomes (HHTTH, HTHTT, TTTTT, etc)
How did we get 32?
Well, there are 2 possible outcomes for the 1st coin flip, 2 possible outcomes for the 2nd coin flip, 2 outcomes for the 3rd flip, 2 outcomes for the 4th flip, and 2 outcomes for the 5th flip,
By the Fundamental Counting Principle (FCP), the TOTAL number of outcomes when flipping 5 coins = (2)(2)(2)(2)(2) = 32

Of course, among those 32 outcomes, there are some outcomes that DO NOT meeting the required condition that we receive at least one head and one tail
There are EXACTLY 2 outcomes that DO NOT meet this required condition.
They are: TTTTT and HHHHH

So, the number of outcomes that DO meet the required condition = 32 - 2 = 30

So, P(child gets at least one head and one tail) = 30/32= 15/16

Answer: C

Cheers,
Brent
Brent Hanneson - Creator of GMATPrepNow.com
Image
Join the discussion

BTGmoderatorDC wrote: ↑
Mon Apr 06, 2020 5:35 pm
If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one tail?

A) 3/4
B) 11/12
C) 15/16
D) 31/32
E) 63/64


OA C

Source: Veritas Prep
The only way to not receive at least one of each would be to have all five of the same kind.

That probability is \(\dfrac{1}{2} \cdot 5\), or \(\dfrac{1}{32}\).

Because there are two ways to get all five of one kind (all heads or all tails), multiply that by \(2\) to get \(\dfrac{1}{16}\).

All of the other possibilities will result in at least one head and one tail, so the correct answer is \(1 - \dfrac{1}{16} = \dfrac{15}{16}\).

Therefore, C
Join the discussion

BTGmoderatorDC wrote: ↑
Mon Apr 06, 2020 5:35 pm
If a child flips a coin five times in a row, what is the probability that she will receive at least one head and one tail?

A) 3/4
B) 11/12
C) 15/16
D) 31/32
E) 63/64


OA C

Source: Veritas Prep
Since each coin has two faces, head and tail, there are 2^5 = 32 different combinations when flipping a coin five times in a row. Of these 32 combinations, only two of them do not have at least one head and one tail. It occurs when all 5 flips turn out to be heads (HHHHH) or all 5 turn out to be tails (TTTTT). The other 30 combinations will have at least one head and one tail, so the probability is 30/32 = 15/16.

Answer: C

Scott Woodbury-Stewart
Founder and CEO
[email protected]

Image

See why Target Test Prep is rated 5 out of 5 stars on BEAT the GMAT. Read our reviews

ImageImage
Join the discussion